OpenStudy (tylerballardwildcat):

What is the probability of rolling an even number on a six sided die and then rolling a 4? >>will medal

OpenStudy (mathgeek289):

Okay. How Man Even Numbers Are There In 1-6?

OpenStudy (mathgeek289):

2, 4, and 6

OpenStudy (mathgeek289):

So, We Have The Probability being 3/6 so far

OpenStudy (mathgeek289):

We Also Know That There Is 1 Four On The Die

OpenStudy (mathgeek289):

So That Is 1/6

OpenStudy (mathgeek289):

What Do We Do Next?

OpenStudy (tylerballardwildcat):

i dont know

OpenStudy (welshfella):

OK You have 2 events First one has probability 3/6 or 1/2 and the second has probability of 1/6. How do we calculate the probabilty of both these 2 events happening. What do we do to the 1/2 abd 1/6?

OpenStudy (welshfella):

* and 1/6

OpenStudy (tylerballardwildcat):

divide

OpenStudy (welshfella):

No Hint the probability of both happening is a lot less than only one happening.

OpenStudy (tylerballardwildcat):

muliply

OpenStudy (mathgeek289):

Yes 1/6 * 3/6 = ?

OpenStudy (mathgeek289):

1 3 -- * -- 6 6

OpenStudy (mathgeek289):

1 * 3 = 3 6 * 6 = 36 So, It Is 3/36 Probability. Simplify. 3 1 -- = --- 36 12